# Segment then Splat: A Unified Approach for 3D Open-Vocabulary Segmentation based on Gaussian Splatting

### v1

@@ Abstract
Open-vocabulary querying in 3D space is crucial for enabling more intelligent perception in applications such as robotics, autonomous systems, and augmented reality. However, most existing methods rely on 2D pixel-level parsing, leading to multi-view inconsistencies and poor 3D object retrieval. Moreover, they are limited to static scenes and struggle with dynamic scenes due to the complexities of motion modeling. In this paper, we propose Segment then Splat, a 3D-aware open vocabulary segmentation approach for both static and dynamic scenes based on Gaussian Splatting. Segment then Splat reverses the long established approach of “segmentation after reconstruction” by dividing Gaussians into distinct object sets before reconstruction. Once the reconstruction is complete, the scene is naturally segmented into individual objects, achieving true 3D segmentation. This approach not only eliminates Gaussian-object misalignment issues in dynamic scenes but also accelerates the optimization process, as it eliminates the need for learning a separate language field. After optimization, a CLIP embedding is assigned to each object to enable open-vocabulary querying. Extensive experiments on various datasets demonstrate the effectiveness of our proposed method in both static and dynamic scenarios.
Figure 1: Comparison of our Segment-then-Splat pipeline and the traditional Splat-then-Segment approach. (a) The traditional Splat-then-Segment pipeline learns a language field alongside the reconstruction of the entire scene. During object queries, it renders Gaussian language embeddings into a 2D feature map to identify relevant pixels based on the input text embedding. (b) In contrast, our Segment-then-Splat pipeline first initialize Gaussians into object-specific sets before reconstruction, ensuring a more precise object-Gaussian correspondence and improving segmentation accuracy.
@@ 1 Introduction
3D open vocabulary querying marks a pivotal step in language-driven interaction with 3D environments, removing the need for predefined labels. This capability is vital for large-scale scene exploration, understanding and editing [3, 29, 30, 16], robotic navigation [7, 19] and manipulation [23, 27], where free-form text bridges human language and machine perception.
@@
To solve this issue, most existing works [9, 21, 28, 36] incorporate a separate language field alongside Gaussian Splatting reconstruction. By rendering the language field into 2D feature maps, they enable pixel-based querying by retrieving relevant pixels based on the input text embedding. However, this approach essentially performs segmentation in 2D space rather than partitioning Gaussians in 3D space, leading to several drawbacks: 1) Inconsistent 2D segmentation results across different views, leading to inaccurate object boundaries. 2) Failure to learn the actual 3D information of objects, making object extraction in 3D challenging and limiting downstream tasks like robot navigation and 3D manipulation. 3) Inapplicability to dynamic scenes, since Gaussians may have varying semantic meanings at different time steps, preventing straightforward extensions to time-varying or moving objects.
@@
Despite the differences in segmentation strategies, whether pixel-based or 3D-based, all existing approaches follow a “reconstruction-then-segmentation” (i.e. splat then segment) paradigm. This approach inherently results in imprecise object boundaries, as each Gaussian may encode geometric and semantic information from multiple objects, leading to segmentation ambiguity.
In this paper, we propose Segment then Splat, a unified framework for 3D-aware open-vocabulary segmentation that can be applied to both static and dynamic scenes. Unlike existing methods that adopt a “splat then segment” approach, our method reverses the process by first initializing each object with a specific set of Gaussians, as shown in Fig. 1. During training, each set of Gaussians is assigned a unique object ID and contributes only to its corresponding object, guided by 2D multi-view mask supervision. By doing so, each Gaussian is dedicated to a single object and thus learns more accurate object geometry. Moreover, since the Gaussian-object correspondence is strictly maintained, our method can be directly applied to dynamic scenes without the risk of Gaussian-object misalignment (i.e. one Gaussian may represent different objects at different time steps). Finally, Segment then Splat requires only one pass of reconstruction and does not depend on learning an additional feature field, significantly improving efficiency. In summary, our key contributions include:
  • We propose Segment then Splat, a unified framework for 3D-aware open-vocabulary segmentation based on Gaussian Splatting.
  • We design a “segmentation then reconstruction” pipeline that enhances object geometry, ensures compatibility with dynamic scenes, and eliminates the need to learn a separate feature field.
  • We introduce a robust object-tracking module capable of mitigating tracking errors in complex scenes.
  • Extensive experiments on diverse static and dynamic datasets demonstrate the superior performance of our method in 3D object segmentation.
@@ 2 Related Work
To address this limitation, many researches focus on how to better model scene dynamics [33, 18, 34, 31, 2, 14, 17]. 4D Gaussian Splatting [34] directly extends 3DGS into four dimensions to facilitate dynamic scene reconstruction. Meanwhile, Deformable-3D-Gaussians [33] leverages a multi-layer perceptron (MLP) to learn per-timestamp positions, rotations, and scales for each Gaussian, effectively capturing object motion and deformation over time. 4DGaussians [31] utilizes multi-resolution HexPlanes [2] to decode features for temporal deformation of 3D Gaussians, while STG [14] encodes changes over time via a temporal opacity term and a polynomial function for each Gaussian, yielding a more detailed representation of dynamic scenes.
Despite these advancements, the above methods act solely as scene representations. After reconstruction, they do not support additional interaction or provide information beyond geometry and texture. In contrast, our approach integrates CLIP [22] embeddings into Gaussian Splatting, assigning each Gaussian an object-level semantic meaning. This capability enables open-vocabulary segmentation, allowing users to retrieve, organize, and query objects within the reconstructed scene through natural language prompts.
@@ 2.2 Language Embedded Scene Representation
Various research has focused on integrating language features [9, 35, 28, 37] into 3D scene representations. LERF [9] pioneered this approach by embedding CLIP features in NeRF. Specifically, it extracts pixel-level CLIP embeddings from multi-scale image crops and trains them alongside NeRF to enable open-vocabulary 3D queries. Building on this idea, LEGaussians [28] incorporates uncertainty and semantic feature attributes into each Gaussian, while introducing a quantization strategy to compress high-level language and semantic features. LangSplat [21] employs a scene-wise language autoencoder to learn language features within a scene-specific latent space, and incorporate SAM mask to enable clear object boundaries in rendered feature images. Despite these advancements, all the above methods essentially perform 2D segmentation when conducting open-vocabulary segmentation, as they compare rendered 2D language features with the input text query embedding.
OpenGaussian [32] leverages contrastive learning to assign a feature embedding to each Gaussian, then applies $K$-means clustering to group Gaussians into multiple object clusters, thereby realizing 3D segmentation. Similarly, GaussianCut [6] utilizes graphcut [1, 4, 5] to segment Gaussians into foreground and background based on user input.
However, all of the above-mentioned methods adhere to a “splat then segment” (i.e. reconstruction then segmentation) pipeline, which inherently results in imprecise object boundaries, since each Gaussian may contain geometry and semantic information from different objects. Furthermore, these methods struggle with dynamic scenes due to Gaussian-object misalignment, preventing their direct application to non-static environments. Dynamic 3D Gaussian Distillation (DGD) [11] distills the feature from LSeg [12] into a feature field to achieve open-vocabulary segmentation, however it still suffers from object-Gaussian misalignment issue as it does not explicitly handle it.
@@ 3 Method
Figure 2: A detailed demonstration of our proposed Segment then Splat pipeline. Our approach first extracts multi-view masks for each object through a robust tracking module, then object IDs are assigned to each initial Gaussian based on these masks, forming distinct object-specific sets. During optimization, object specific loss $\mathcal{L}_{\text{obj}}$ is used to enforce Gaussian-object correspondence and thus resulting in more accurate object geometries. Finally, we associate each Gaussian group with CLIP embeddings, enabling open-vocabulary queries.
We introduce Segment then Splat, a unified approach for 3D open-vocabulary segmentation based on Gaussian Splatting, as illustrated in Fig. 2. The process begins by extracting multi-view masks for each object through a robust object tracking module (Sec. 3.2), ensuring reliable detection and mask generation. Next, each Gaussian initialized by COLMAP [26, 25] is assigned an object ID according to these masks, partitioning the entire scene into multiple object-specific Gaussian sets (Sec. 3.3). During optimization & reconstruction (Sec. 3.4), each Gaussian contributes exclusively to its assigned object, preserving Gaussian-object correspondence and resulting in more accurate object geometries. After reconstruction, CLIP embeddings are associated with each group of Gaussians (Sec. 3.5), enabling open-vocabulary queries. Besides, three levels of granularity (large, middle, and small) are introduced to facilitate object retrieval at different scales.
@@ 3.1 Preliminary: 3D and 4D Gaussian Splatting
3DGS. 3D Gaussian Splatting represents a scene using a collection of 3D ellipsoids, each modeled as an anisotropic 3D Gaussian. Each Gaussian is parameterized by a mean $x$, which defines the center of the ellipsoid, a covariance matrix $\Sigma$ , which determines its shape, as shown in Eq. 1. The color of the Gaussian is defined using spherical harmonics.
$G(x)=e^{-\frac{1}{2}(x)^{T}\Sigma^{-1}(x)}$ (1)
@@ 3.2 Robust Object Tracking
Given a set of input images $\{I_{i}\}_{i=0}^{n}$, our goal is to extract multi-view masks for all objects at different granularity levels (i.e. large, middle and small) in the scene. We begin by leveraging Segment Anything (SAM) [10] with grid-based point prompting to obtain initial static object masks of different granularities in the first input frame $I_{0}$. Subsequently, SAM2 [24] is employed to track objects throughout the sequence based on the extracted mask from $I_{0}$. However, this process presents several challenges: 1) Some objects may not appear in the first frame, leading to their exclusion from tracking. 2) Due to grid-based prompting, one pixel may be tracked multiple times into different masks, either belonging to a single object or a part of an object. 3) If an object temporarily disappears and reappears later in the scene, tracking may be lost. To address these issues, we propose several post-processing strategies.
Detect Any New Objects. To ensure newly appearing objects are captured, we introduce a detection mechanism at fixed intervals of $\Delta t$. Specifically, we compare the ratio of the segmented region in frame $I_{t+\Delta t}$ to that in frame $I_{t}$. A significant decline in this ratio indicates the potential presence of new objects. At this point, we re-segment the scene and analyze the intersection between the new and previous segmentation results. Objects with minimal overlap with prior masks are identified as new objects and added to the static segmentation results. SAM2 then continues tracking based on this updated segmentation.
@@
Handling Lost Tracking. When an object’s tracking is lost, it may be incorrectly treated as a new object upon its reappearance in subsequent frames, leading to multiple instances representing the same object. In these scenarios, we resolve this issue using the approach described in Sec. 3.3.
@@ 3.3 Object-Specific Gaussian Initialization
As our method follows a “segmentation then reconstruction” strategy, we first segment the Gaussians initialized by COLMAP into distinct sets, each representing a different object. Each Gaussian is assigned three object IDs, corresponding to three granularity levels. To determine these IDs, we analyze the visibility of each Gaussian center across all views and identify the corresponding object mask region in which it resides. After object IDs are determined, we handle the lost tracking issue stated in Sec. 3.2. When an object is tracked multiple times as different instances, the corresponding Gaussians should share a similar geometric center and appearance (e.g. color). Therefore, we refine the segmentation by merging Gaussians and its corresponding object mask, if they exhibit a closely aligned “geometric-appearances distance”, defined as follows:
@@ 3.4 Optimization & Reconstruction
$\mathcal{L}_{\text{obj}}=\mathcal{L}_{1}(M_{i}^{p}\otimes I_{i},\hat{I_{i}^{p}}),$ (6)
where $\hat{I}_{i}$ is the rendered $i^{\text{th}}$ image, and $\hat{I}_{i}^{p}$ and $M_{i}^{p}$ represent the rendered $p^{\text{th}}$ object in the $i^{\text{th}}$ image and its corresponding mask, respectively, as extracted in Sec. 3.2. The overall loss function is as follows:
@@
Throughout the entire reconstruction process, all densification and cloning operations are performed strictly within each object-specific Gaussian set, preserving the Gaussian-object correspondence. Additionally, we implement a Gaussian persistence mechanism to ensure that each set of object-specific Gaussians is not completely pruned, thereby preventing scenarios where an object could lose all its associated Gaussians.
@@
Optimization on Multiple Granularities. Noticed that we have three level granularities: large, middle and small. To ensure effective optimization, we must account for these varying scales when determining the optimization order. Since smaller objects are always part of larger ones, they should be optimized first. If the order is reversed, i.e. optimizing smaller objects after larger ones, the internal structure of the larger objects may become disorganized again, as shown in Fig. 3. Thus the $\mathcal{L}_{\text{obj}}$ will finally be formulated as:
$\mathcal{L}_{\text{obj}}=\left\{\begin{array}[]{ll}\mathcal{L}_{\text{obj\_S}},&\text{stage1}\\ \mathcal{L}_{\text{obj\_S}}+\mathcal{L}_{\text{obj\_M}},&\text{stage2}\\ \mathcal{L}_{\text{obj\_S}}+\mathcal{L}_{\text{obj\_M}}+\mathcal{L}_{\text{obj\_L}},&\text{stage3}\end{array},\right.$ (8)
Figure 3: A demonstration of how the optimization order of different granularity will affect the reconstruction result. When small-level objects are optimized first, the structures of both small-level and middle-level objects are well-preserved. In contrast, optimizing middle-level objects first results in well-maintained middle-level structures but chaotic internal structures at the small level due to the absence of supervision.
@@
Partial Mask Filtering. In 3D-aware segmentation, objects can be reconstructed even from views where they would typically appear occluded. Consequently, multi-view 2D masks provided as supervision can introduce discrepancies, as they do not account for occluded regions revealed during rendering, as shown in Fig. 5. This discrepancy can lead to incorrect constraints, ultimately distorting the geometric structure of the 3D objects. The original Gaussian Splatting (3DGS) framework can partially resolve this issue by leveraging multi-view consistency, but errors from unreliable masks persists. To robustly address this, we propose a partial mask filtering strategy applied at the end of training. Specifically, we render each reconstructed object into 2D images, compute their Intersection-over-Union (IoU) against the provided masks, and discard masks exhibiting low IoU scores. This ensures that only consistent, accurate masks inform the final optimization, significantly enhancing the geometric fidelity of the reconstructed 3D objects.

### v2

# Segment then Splat: Unified 3D Open-Vocabulary Segmentation via Gaussian Splatting

@@ Abstract
Open-vocabulary querying in 3D space is crucial for enabling more intelligent perception in applications such as robotics, autonomous systems, and augmented reality. However, most existing methods rely on 2D pixel-level parsing, leading to multi-view inconsistencies and poor 3D object retrieval. Moreover, they are limited to static scenes and struggle with dynamic scenes due to the complexities of motion modeling. In this paper, we propose Segment then Splat, a 3D-aware open vocabulary segmentation approach for both static and dynamic scenes based on Gaussian Splatting. Segment then Splat reverses the long established approach of “segmentation after reconstruction” by dividing Gaussians into distinct object sets before reconstruction. Once reconstruction is complete, the scene is naturally segmented into individual objects, achieving true 3D segmentation. This design eliminates both geometric and semantic ambiguities, as well as Gaussian–object misalignment issues in dynamic scenes. It also accelerates the optimization process, as it eliminates the need for learning a separate language field. After optimization, a CLIP embedding is assigned to each object to enable open-vocabulary querying. Extensive experiments one various datasets demonstrate the effectiveness of our proposed method in both static and dynamic scenarios.
@@ 1 Introduction
Figure 1: Traditional 3D Open-Vocabulary Segmentation vs. our Segment-then-Splat Pipeline. (a) The traditional Splat-then-Segment pipeline learns a language field alongside the reconstruction of the entire scene. During object queries, it renders Gaussian language embeddings into a 2D feature map to identify relevant pixels based on the input text embedding. (b) In contrast, our Segment-then-Splat pipeline first initialize Gaussians into object-specific sets before reconstruction, ensuring a more precise object-Gaussian correspondence and improving segmentation accuracy.
3D open-vocabulary querying marks a pivotal step in language-driven interaction with 3D environments, removing the need for predefined labels. This capability is vital for large-scale scene exploration, scene understanding, robotic navigation [1, 2, 3] and manipulation [4, 5, 6, 7], where free-form text bridges human language and machine perception.
@@
To solve this issue, most existing works [9, 10, 11, 12] incorporate a separate language field alongside Gaussian Splatting reconstruction. By rendering the language field into 2D feature maps, they enable pixel-based querying by retrieving relevant pixels based on the input text embedding. However, this approach essentially performs segmentation in 2D space rather than partitioning Gaussians in 3D space, leading to several drawbacks: 1) Inconsistent 2D segmentation across different views, leading to inaccurate object boundaries. 2) Failure to capture true 3D object information, complicating 3D object extraction and limiting downstream tasks like robot navigation and 3D manipulation. 3) Inapplicability to dynamic scenes, since Gaussians may have varying semantic meanings at different time steps, preventing straightforward extensions to time-varying or moving objects.
@@
Despite the differences in segmentation strategies, whether pixel-based or 3D-based, all existing approaches follow a “reconstruction then segmentation” (i.e., splat then segment) paradigm. This approach inherently results in imprecise object boundaries, as each Gaussian may encode geometric and semantic information from multiple objects, leading to geometric and semantic ambiguity.
In this paper, we propose Segment then Splat, a unified framework for 3D-aware open-vocabulary segmentation that can be applied to both static and dynamic scenes. Unlike existing methods that adopt a “splat then segment” approach, our method reverses the process by first initializing each object with a specific set of Gaussians, as shown in Fig.˜1. During training, each set of Gaussians is assigned a unique object ID and contributes only to its corresponding object, guided by 2D multi-view mask supervision. By doing so, each Gaussian is dedicated to a single object and thus learns more accurate object geometry. Moreover, since the Gaussian-object correspondence is strictly maintained, our method can be directly applied to dynamic scenes without the risk of Gaussian-object misalignment (i.e., one Gaussian may represent different objects at different time steps). Finally, Segment then Splat requires only one pass of reconstruction and does not depend on learning an additional feature field, significantly improving efficiency. In summary, our key contributions include:
  • We propose Segment then Splat, a novel paradigm that segments Gaussians into object sets before reconstruction. This enables unified static/dynamic open-vocabulary segmentation, eliminates auxiliary language fields, and significantly reduces training complexity.
  • Our framework features a robust object tracking module that maintains spatial-temporal consistency of object-specific Gaussians in the scene, ensuring accurate segmentation and motion modeling while preventing misalignment.
  • Learning object-specific Gaussians from the outset preserves explicit object-Gaussian correspondence, eliminating geometric and semantic ambiguity, yielding superior 3D geometries and multi-level segmentation granularity.
  • Extensive experiments demonstrate state-of-the-art performance across diverse static and dynamic datasets in 3D open-vocabulary segmentation, object geometry accuracy, and computational efficiency.
@@ 2 Related Work
To address this limitation, Dynamic 3D Gaussians [15] employs a table-based strategy, storing each Gaussian’s mean and variance at every timestamp. 4D Gaussian Splatting [16] extends 3DGS into four dimensions, adding a temporal component to facilitate dynamic scene reconstruction. Meanwhile, Deformable 3D Gaussians [17] leverages a multi-layer perceptron (MLP) to learn per-timestamp positions, rotations, and scales for each Gaussian, effectively capturing object motion and deformation over time. 4DGaussians [18] utilizes multi-resolution HexPlanes [19] to decode features for temporal deformation of 3D Gaussians, while STG [20] uses a temporal opacity term and a polynomial function for each Gaussian, yielding a more detailed representation of dynamic scenes.
Despite these advancements, the above methods act solely as scene representations. After reconstruction, they do not support additional interaction or provide information beyond geometry and texture. In contrast, our approach enhances interaction by integrating CLIP [21] embeddings into Gaussian Splatting. This assigns semantic meaning to each Gaussian, enabling open-vocabulary segmentation where users can retrieve, organize, and query objects using natural language prompts.
@@ 2.2 Language Embedded Scene Representation
Various research has focused on integrating language features [9, 22, 11, 23, 24] into 3D scene representations. LERF [9] pioneered this approach by embedding CLIP features in NeRF. Specifically, it extracts pixel-level CLIP embeddings from multi-scale image crops and trains them alongside NeRF to enable open-vocabulary 3D queries. Building on this idea, LEGaussians [11] incorporates uncertainty and semantic feature attributes into each Gaussian, while introducing a quantization strategy to compress high-level language and semantic features. LangSplat [10] employs a scene-wise language autoencoder to learn language features within a scene-specific latent space, and incorporate SAM mask to enable clear object boundaries in rendered feature images. Despite these advancements, all the above methods essentially perform 2D segmentation when conducting open-vocabulary segmentation, as they compare rendered 2D language features with the input text query embedding. OpenGaussian [13] leverages contrastive learning to assign a feature embedding to each Gaussian, then applies $K$-means clustering to group Gaussians into multiple object clusters, thereby realizing 3D segmentation. Similarly, GaussianCut [14] utilizes graphcut [25, 26, 27] to segment Gaussians into foreground and background based on user input.
However, all of the above-mentioned methods adhere to a “splat then segment” (i.e., reconstruction then segmentation) pipeline, which inherently results in imprecise object boundaries, since each Gaussian may contain geometry and semantic information from different objects. Furthermore, these methods struggle with dynamic scenes due to Gaussian-object misalignment, preventing their direct application to non-static environments. Dynamic 3D Gaussian Distillation (DGD) [28] distills the feature from LSeg [29] into a feature field to achieve open-vocabulary segmentation. 4D LangSplat [30] further leverages Multimodal Large Language Models (MLLMs) through object-wise video prompting to enhance both time-sensitive and time-agnostic open-vocabulary understanding. Similarly, 4-LEGS [31] distills spatio-temporal language features into 4DGS for event localization from text prompts. Nevertheless, these methods require specific modifications for dynamic scenarios and still suffer from object–Gaussian misalignment.
@@ 3 Method
Figure 2: Demonstration of Segment then Splat pipeline. We first extracts multi-view masks for each object through a robust tracking module, then object IDs are assigned to each initial Gaussian based on these masks, forming distinct object-specific sets. During optimization, object specific loss $\mathcal{L}_{\text{obj}}$ is used to enforce Gaussian-object correspondence and thus resulting in more accurate object geometries. Finally, a CLIP embedding is assigned to each Gaussian group for open-vocabulary queries.
We introduce Segment then Splat, a unified approach for 3D open-vocabulary segmentation based on Gaussian Splatting, as illustrated in Fig.˜2. The process begins by extracting multi-view masks for each object through a robust object tracking module (Sec.˜3.2), ensuring reliable detection and mask generation. Next, each Gaussian initialized by COLMAP [32, 33] is assigned an object ID according to these masks, partitioning the entire scene into multiple object-specific Gaussian sets (Sec.˜3.3). During optimization & reconstruction (Sec.˜3.4), each Gaussian contributes exclusively to its assigned object, preserving Gaussian-object correspondence and resulting in more accurate object geometries. After reconstruction, CLIP embeddings are associated with each group of Gaussians (Sec.˜3.5), enabling open-vocabulary queries. Besides, three levels of granularity (large, middle, and small) are introduced to facilitate object retrieval at different scales.
@@ 3.1 Preliminary: 3D and 4D Gaussian Splatting
3DGS. 3D Gaussian Splatting represents a scene using a collection of 3D ellipsoids, each modeled as an anisotropic 3D Gaussian. Each Gaussian is parameterized by a mean $x$, which defines the center of the ellipsoid, a covariance matrix $\Sigma$ , which determines its shape, as shown in Eq.˜1. The color of the Gaussian is defined using spherical harmonics.
$G(x)=e^{-\frac{1}{2}(x)^{T}\Sigma^{-1}(x)}.$ (1)
@@ 3.2 Robust Object Tracking
Given a set of input images $\{I_{i}\}_{i=0}^{n}$, our goal is to extract multi-view masks for all objects at different granularity levels (i.e., large, middle and small) in the scene. We begin by leveraging Segment Anything (SAM) [34] with grid-based point prompting to obtain initial static object masks of different granularities in the first input frame $I_{0}$. Subsequently, SAM2 [35] is employed to track objects throughout the sequence based on the extracted mask from $I_{0}$. However, this process presents several challenges: 1) Some objects may not appear in the first frame, leading to their exclusion from tracking. 2) Due to grid-based prompting, one pixel may be tracked multiple times into different masks, either belonging to a single object or a part of an object. 3) If an object temporarily disappears and reappears later in the scene, tracking may be lost. To overcome these issues and improve the robustness of object tracking, we propose three targeted post-processing strategies that dynamically detect new objects, resolve mask conflicts, and recover from tracking failures. These strategies transform SAM-based segmentation into a flexible, scene-adaptive tracking pipeline, which serves as the critical foundation for subsequent steps in our methodology.
Detect Any New Objects. To capture newly appearing objects, we introduce a detection mechanism at fixed intervals of $\Delta t$. Specifically, we compare the segmented region ratio between frames $I_{t+\Delta t}$ and $I_{t}$. A significant decline in this ratio indicates the potential presence of new objects. At this point, we re-segment the scene and analyze the intersection between the new and previous segmentation results. Objects with minimal overlap with prior masks are identified as new objects and added to the static segmentation results. SAM2 then continues tracking based on this updated segmentation.
@@
Handling Lost Tracking. When an object’s tracking is lost, it may be incorrectly treated as a new object upon its reappearance in subsequent frames, leading to multiple instances representing the same object. In these scenarios, we resolve this issue using the approach described in Sec.˜3.3.
@@ 3.3 Object-Specific Gaussian Initialization
As our method follows a “segmentation then reconstruction” strategy, we first segment the Gaussians initialized by COLMAP into distinct sets, each representing a different object. Each Gaussian is assigned three object IDs, corresponding to three granularity levels. To determine these IDs, we analyze the visibility of each Gaussian center across all views and identify the corresponding object mask region in which it resides. After object IDs are determined, we handle the lost tracking issue stated in Sec.˜3.2. When an object is tracked multiple times as different instances, the corresponding Gaussians should share a similar geometric center and appearance (e.g. color). Therefore, we refine the segmentation by merging Gaussians and its corresponding object mask, if they exhibit a closely aligned “geometric-appearances distance”, defined as follows:
@@ 3.4 Optimization & Reconstruction
[…]
where $\hat{I}_{i}$ is the rendered $i^{\text{th}}$ image, and $\hat{I}_{i}^{p}$ and $M_{i}^{p}$ represent the rendered $p^{\text{th}}$ object in the $i^{\text{th}}$ image and its corresponding mask, respectively, as described in Sec.˜3.2. $\otimes$ denotes element-wise multiplication. The overall loss function is as follows:
@@
Throughout the entire reconstruction process, all densification and cloning operations are performed strictly within each object-specific Gaussian set, preserving the Gaussian-object correspondence. Besides, a Gaussian persistence mechanism is designed to ensure that each set of Gaussians is not completely pruned, thereby preventing scenarios where an object could lose all its associated Gaussians.
@@
Optimization on Multiple Granularities. Noticed that we have three level granularities: large, middle and small. To ensure effective optimization, we must account for these varying scales when determining the optimization order. Since smaller objects are always part of larger ones, they should be optimized first. If the order is reversed, i.e., optimizing smaller objects after larger ones, the internal structure of the larger objects may become disorganized again, as shown in Fig.˜3. Thus the $\mathcal{L}_{\text{obj}}$ will finally be formulated as:
$\mathcal{L}_{\text{obj}}=\begin{cases}\mathcal{L}_{\text{obj\_S}},&\text{stage1}\\ \mathcal{L}_{\text{obj\_S}}+\mathcal{L}_{\text{obj\_M}},&\text{stage2}\\ \mathcal{L}_{\text{obj\_S}}+\mathcal{L}_{\text{obj\_M}}+\mathcal{L}_{\text{obj\_L}},&\text{stage3}\end{cases}$ (8)
@@
Figure 3: A demonstration of how the optimization order affects reconstruction. Optimizing small-level objects first preserves both small- and middle-level structures, while starting with middle-level ones leads to well-maintained middle-level but chaotic small-level regions due to lack of supervision.
Partial Mask Filtering. In 3D-aware segmentation, objects can be reconstructed even from views where they would typically appear occluded. Consequently, multi-view 2D masks provided as supervision can introduce discrepancies, as they do not account for occluded regions revealed during rendering, as shown in Fig.˜5. This discrepancy can lead to incorrect constraints, ultimately distorting the geometric structure of the 3D objects. The original 3DGS framework can partially resolve this issue by leveraging multi-view consistency, but errors from unreliable masks persists. To robustly address this, we propose a partial mask filtering strategy applied at the end of training. Specifically, we render each reconstructed object into 2D images, compute their Intersection-over-Union (IoU) against the provided masks, and discard masks exhibiting low IoU scores. This ensures that only consistent, accurate masks inform the final optimization, significantly enhancing the geometric fidelity of the reconstructed 3D objects.